This question was on testprepreview.com which I was using to practice for the COMPASS test. My answer is different from what they say the correct answer is. Could you please show me the step by step solution?
"If Sally can paint a house in 4 hours, and John can paint the same house in 6 hours, how long will it take both of them to paint the house together?" I figured Sally could paint 1/4 of the house in 1 hour and John could paint 1/6 of the house in 1 hour. So I added them together to get they could paint 5/12 of the house in 1 hour. I figured if they could paint 5/12 of the house in 1 hour then to paint 12/12 of the house, the whole thing would take 2 1/6 hours, which is 2 hours and 10 minutes. They gave the correct answer as 2 hours and 24 minutes. I need to understand how they worked this. Thank you.
Joyce M. Smith
( a new college student at age 62)

Answer by checkley77(12844)

This is one of the simple straightforward solutions: s*j/(s+j)
4*6/(4+6)=24/10=2.4 hours to complete the job working together.